What interest rate, compounded quarterly, is equivalent to a 9.31% effective interest rate?
1
Expert's answer
2021-09-12T23:08:00-0400

Answer:

 "F=1.09P"

Now,

"P(1+i)^4=1.09P"

"(1+i)^4=1.09"

"(1+i)=1.09^{^1\/4}"

"(1+i)=1.09^{^1\/4}-1"

"=0.02177818086"

the quarterly rate is about"=2.18\%"


RESULT

"2.18\%"